# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

### v1

@@ 3.2 MOSFiT modeling: physical parameters
In almost all cases, we find no obvious differences in the parameter distributions between any two of the three radio classes (see Table 8 in Appendix C for the full results). The one exception is the photospheric radius normalization parameter $R_{\rm ph,0}$, which is larger for the delayed radio TDEs than for the prompt radio TDEs ($p=0.026$). There is also a marginal suggestion ($p\sim 0.05$) that the TDEs with delayed radio emission have lower SMBH masses than either the TDEs with prompt radio emission or than the host+none group – but only when $M_{\rm BH,host}$ is used; no significant differences are found between the $M_{\rm BH,mosfit}$ distributions for these groups. We plot the distributions of $R_{\rm ph,0}$, $M_{\rm BH,host}$, and $M_{\rm BH,mosfit}$ in Figure 1. While achieving three $p$ values $\lesssim 0.05$ out of 36 statistical tests performed is not a highly significant result, in Section 4 we briefly explore the possible implications if these correlations are verified with a larger TDE sample in future work. We emphasize that null results in our current sample do not necessarily imply that correlations between the radio properties and the various parameters tested are not present, given the limitations of existing data. For example, several of the TDEs that we classify as radio type Delayed have no radio observations at $t<6$ months, meaning that we have no way of knowing if they also displayed prompt radio emission and are thus misclassified here, potentially diluting any underlying trends (Section 4.2.3).
@@
Figure 3: Same as Figure 2 for the remaining TDEs in our sample with a delayed radio emission episode and converged MOSFiT models. For six of these TDEs, C24 provide a rough estimate of the outflow launch time (dashed vertical red lines). For AT2019dsg, the dashed line indicates the possible launch time of the second radio flare. For PS16dtm, the open circle denotes a marginal possible detection discovered in our reanalysis of the data (Section 2.1). C24 do not estimate a launch time for AT2020neh due to its faintness.
@@
3.3 MOSFiT modeling: accretion rate
@@
In Figures 2-4 we show the $\dot{M}$ curves together with the radio and X-ray light curves for all TDEs from our sample with unambiguously transient radio emission and converged MOSFiT models666i.e., all TDEs with radio type “Prompt” or “Delayed” in Table 1 except for AT2018fyk, whose optical emission cannot be well-fit by MOSFiT. Further late-time radio monitoring of the TDEs with radio types Host or None is ongoing and will be presented in future work.. Specifically, Figure 2 shows TDEs for which we have a reasonably well-constrained estimate of the outflow launch time [^7] ASASSN-14ae, AT2018hco, AT2018hyz, and AT2019eve have multiple radio SEDs at late times, so it is possible to use the evolution of the derived physical size of the radio-emitting region to obtain a rough estimate of the time at which their delayed outflows were launched (Cendes et al. 2022b and C24). Similarly, Horesh et al. (2021b) model the multi-frequency radio light curves of iPTF16fnl to constrain the likely radio outflow launch time to $\sim 56-153$ days post-discovery. Multi-wavelength observations of ASASSN-15oi suggest that an outflow launched at the time of the soft X-ray peak may provide a good explanation for its second radio flare (Horesh et al., 2021a; Hajela et al., 2024)., Figure 3 shows TDEs with delayed radio emission where we have at most an approximate constraint on the outflow launch time from C24888plus AT2019dsg, for which we show the approximate onset time of its secondary delayed peak from C24, and AT2020neh, which is too faint for an accurate launch estimate., and Figure 4 shows the remaining TDEs with prompt radio emission. We normalize $\dot{M}$ to $\dot{M}_{\rm Edd}$, where $\dot{M}_{\rm Edd}$ is the accretion rate required to power emission equal to the Eddington luminosity of the SMBH assuming a radiative efficiency $\epsilon$; $L_{\rm Edd}=\epsilon\dot{M}_{\rm Edd}c^{2}$. We use the radiative efficiency from MOSFiT for each TDE and our two independent estimates of the SMBH mass from Section 3.1. We find that most TDEs in our sample achieve a peak $\dot{M}$ that approaches or mildly exceeds $\dot{M}_{\rm Edd}$. The peak radio luminosity of a TDE does not appear to correlate with the maximum $\dot{M}/\dot{M}_{\rm Edd}$ achieved onto its SMBH.
@@
We next investigate if the onset time of the radio emission in our TDEs correlates with any specific $\dot{M}/\dot{M_{\rm Edd}}$ ratio, beginning with the TDEs in Figure 2 where we have the best constraints. We find that three of these TDEs (AT2018hco, ASASSN-15oi, and iPTF16fnl) have outflows that were plausibly launched when the accretion rate was a few percent of $\dot{M}_{\rm Edd}$, although for the latter two objects we are limited by the systematic uncertainties on the SMBH masses. ASASSN-15oi’s and iPTF16fnl’s radio emission may also have been launched earlier, possibly when $\dot{M}$ was still super-Eddington. [^9] This is particularly likely for ASASSN-15oi, whose multi-wavelength observations suggest that accretion in this TDE may be delayed a few hundred days relative to our MOSFiT prediction (Hajela et al., 2024). AT2018hyz and AT2019eve’s radio outflows do not appear until several hundred days after the $0.03\dot{M}/\dot{M}_{\rm Edd}$ transition, implying that either accretion is delayed relative to the MOSFiT prediction for these two events, or that some other process is responsible for their radio brightenings. [^10] The launch time from Cendes et al. (2022b) assumes that AT2018hyz’s radio outflow is non-relativistic; if its radio emission is instead due to a relativistic off-axis jet, then a much earlier launch time coinciding with the predicted period of super-Eddington accretion is possible (Matsumoto & Piran, 2023; Sfaradi et al., 2024). Finally, ASASSN-14ae’s very late outflow launch time is difficult to reconcile with either state transition, assuming the outflow launch time and the mass of its SMBH are both correct. This may suggest that either accretion disk formation is significantly delayed in this TDE relative to the prediction from MOSFiT, or some other unrelated process is required to explain its radio rise.
@@
Although the onset time of radio emission is less clear for the TDEs in Figure 3, we can roughly approximate the launch time of each outflow based on the radio light curve (C24; dashed red lines in Figure 3). [^11] We note that our new tentative detection of PS16dtm (Section 2.1; open circle in Figure 3) pre-dates the launch time suggested by C24 for this TDE; it is therefore likely that its radio outflow was launched significantly earlier ($\sim 1$ year post-discovery). The systematic uncertainties on the SMBH masses and, in some cases, the sparse temporal sampling of the radio light curves preclude a definitive association of the radio emission with a particular accretion rate for these TDEs. It is nevertheless allowed within the uncertainties that the delayed radio emission in all of these events (including the secondary peak in the light curve of AT2019dsg) is driven by a state change in the accretion flow associated with $\dot{M}$ declining to $\lesssim 0.03\dot{M}_{\rm Edd}$. Again, in most cases, peak accretion onto the SMBH would have to be delayed by several hundred days or more relative to the MOSFiT predictions to associate the outflow launch times with super-Eddington accretion instead. [^12] PS16dtm and AT2019ehz are the two possible exceptions, if their host galaxy-derived SMBH masses rather than their MOSFiT-derived SMBH masses are assumed to be correct.
@@
3.4 X-ray light curves and the Fundamental Plane
@@
The X-ray properties of our TDEs may also connect to the instantaneous accretion rate onto the SMBH (e.g., Metzger 2022). Together with our SMBH mass measurements (Section 3.1), we can use our radio and X-ray light curves to track the evolution of individual TDEs in the $L_{X}$-$L_{R}$ phase space over thousands of days and compare the properties of our sample to the Fundamental Plane of black hole activity. The Fundamental Plane is a well-established correlation between radio luminosity, X-ray luminosity, and black hole mass that holds across many orders of magnitude for black holes accreting at low rates (Merloni et al., 2003; Falcke et al., 2004). If delayed radio emission is produced in TDEs after the accretion rate has dropped to similarly low levels, we may expect these TDEs to obey the same relation at late times.
We show our TDE sample on the Fundamental Plane in Figure 5, using the SMBH masses derived from the host galaxy properties [^13] Using $M_{\rm BH,host}$ rather than $M_{\rm BH,mosfit}$ allows us to include DES14C1kia and AT2018fyk, which do not have converged MOSFiT models. (Table 2). The original sample of AGN from Merloni et al. (2003) are also shown for comparison. As this relation is calibrated using radio observations collected at 5 GHz, we restrict ourselves to radio data taken at similar frequencies ($\nu<10$ GHz). We furthermore only show TDEs for which quasi-simultaneous radio and X-ray data are available, i.e., observations for which $\delta t/t<0.25$. For TDEs with high-cadence observations, if multiple radio/X-ray data point pairs fall within this tolerance, we select the data pairs with the smallest separation in time and frequency space, such that each available data point is used at most once in the figure. We highlight six TDEs of particular interest in the figure legend. Among the TDEs with sufficiently good radio and X-ray coverage, we see that all of our prompt radio TDEs and many of our delayed radio TDEs are consistent with the Fundamental Plane. Conversely some TDEs, such as ASASSN-15oi, move away from the Plane at late times.
4 Discussion
We now attempt to synthesize our multi-wavelength analyses into a physical model, or models, to explain delayed radio emission in TDEs. Any model must be able to explain our key observational findings from Section 3, which we briefly summarize here:

### v2

@@ III.2 MOSFiT modeling: physical parameters
We first use these results to search for possible correlations between the radio properties of our TDEs and other physical parameters of the disruption. We divide our sample into three groups for this analysis: those with radio type “Prompt” in Table 1 (i.e., TDEs that show early/prompt radio emission), “Delayed” TDEs in Table 1 (those that show late/delayed radio emission), and “None” or “Host” TDEs in Table 1 (those that show no radio emission or only host or faint ambiguous radio emission). We also considered the None and Host TDEs as two independent categories, but this does not significantly change our results. We then performed an Anderson-Darling (AD) test [^6] The Anderson-Darling test is similar to the Kolmogorov–Smirnov (KS) test used by some previous authors (e.g., Nicholl et al. 2022), but it has been shown to have added sensitivity to the tails of the distributions in datasets with large numbers of free parameters (Hou et al., 2009). We note that a KS test applied to our data gives similar results. for each of the parameters listed in Table 3 plus $M_{\rm BH,host}$ (the SMBH masses derived from host galaxy properties presented in Table 2). In almost all cases, we find no obvious differences in the parameter distributions between any two of the three radio classes (Table 4). The one exception is the photospheric radius normalization parameter $R_{\rm ph,0}$, which is larger for the delayed radio TDEs than for the prompt radio TDEs ($p=0.026$). There is also a marginal suggestion ($p\sim 0.05$) that the TDEs with delayed radio emission have lower SMBH masses than either the TDEs with prompt radio emission or than the host+none group – but only when $M_{\rm BH,host}$ is used; no significant differences are found between the $M_{\rm BH,mosfit}$ distributions for these groups. We plot the distributions of $R_{\rm ph,0}$, $M_{\rm BH,host}$, and $M_{\rm BH,mosfit}$ in Figure 1. While achieving three $p$ values $\lesssim 0.05$ out of 36 statistical tests performed is not a highly significant result, in Section IV we briefly explore the possible implications if these correlations are verified with a larger TDE sample in future work. We emphasize that null results in our current sample do not necessarily imply that correlations between the radio properties and the various parameters tested are not present, given the limitations of existing data.
[…]
=1.8in {rotatetable}
Table 3: MOSFiT Parameters
PS16dtm $84^{+26}_{-25}$ $0.63^{+0.14}_{-0.19}$ $7.3^{+0.2}_{-0.2}$ $17.3^{+1.3}_{-0.8}$ $4^{+2}_{-1}$ $5^{+0.3}_{-0.1} 0.0$$05^{+0.05}_{-0.02}$ $104^{+7}_{-11}$ $0.2^{+0.2}_{-0.1}$ $2.51^{+0.08}_{-0.08}$
iPTF16fnl $0.07^{+0.51}_{-0.06}$ $1.04^{+0.05}_{-0.57}$ $6.1^{+0.3}_{-0.4}$ $20.8^{+0.1}_{-3.0}$ $3.1^{+1.5}_{-0.8}$ $105^{+1.113}_{-0.008} 0.0$$005^{+0.003}_{-0.001}$ $17^{+4}_{-9}$ $0.4^{+0.8}_{-0.1}$ $0.5^{+0.3}_{-0.6}$
OGLE17aaj $5^{+25}_{-5}$ $0.98^{+0.12}_{-0.16}$ $7.22^{+0.13}_{-0.29}$ $19^{+2}_{-1}$ $0.13^{+0.10}_{-0.03}$ $7^{+1.1}_{-0.4} 0.0$$03^{+0.07}_{-0.02}$ $14^{+3}_{-4}$ $0.5^{+0.9}_{-0.3}$ $6.4^{+3.2}_{-5.0}$
AT2017eqx $0.04^{+0.91}_{-0.04}$ $1.0^{+0.1}_{-0.2}$ $6.7^{+0.3}_{-0.3}$ $21.0^{+0.1}_{-2.9}$ $5^{+26}_{-3}$ $2^{+0.3}_{-0.1} 0.0$$02^{+0.05}_{-0.01}$ $39^{+13}_{-17}$ $0.3^{+0.8}_{-0.2}$ $1.9^{+1.5}_{-2.6}$
AT2018zr $0.01^{+0.31}_{-0.01}$ $0.87^{+0.07}_{-0.07}$ $6.79^{+0.05}_{-0.11}$ $20.92^{+0.05}_{-0.07}$ $9^{+3}_{-2}$ $11^{+0.07}_{-0.04} 0.0$$04^{+0.02}_{-0.02}$ $8.0^{+0.7}_{-0.8}$ $0.2^{+0.1}_{-0.0}$ $1.5^{+0.2}_{-0.3}$
AT2018bsi $0.1^{+3.1}_{-0.1}$ $0.9^{+0.3}_{-0.2}$ $6.9^{+0.3}_{-0.4}$ $20.6^{+0.2}_{-0.8}$ $1.5^{+2.2}_{-0.9}$ $7^{+0.6}_{-0.2} 0.0$$011^{+0.033}_{-0.005}$ $78^{+32}_{-58}$ $0.3^{+0.8}_{-0.2}$ $2.1^{+1.0}_{-2.3}$
AT2018hco $0.1^{+2.4}_{-0.1}$ $0.77^{+0.10}_{-0.10}$ $6.87^{+0.08}_{-0.27}$ $19^{+2}_{-2}$ $1.7^{+0.2}_{-0.2}$ $12^{+0.10}_{-0.05} 0.0$$08^{+0.05}_{-0.04}$ $8^{+1}_{-2}$ $0.32^{+0.27}_{-0.09}$ $2.4^{+0.3}_{-0.5}$
AT2018lna $0.10^{+1.90}_{-0.09}$ $0.97^{+0.12}_{-0.19}$ $6.65^{+0.15}_{-0.15}$ $18^{+2}_{-2}$ $2.0^{+1.1}_{-0.5}$ $6^{+0.3}_{-0.2} 0.0$$007^{+0.003}_{-0.003}$ $15^{+5}_{-7}$ $0.4^{+0.3}_{-0.1}$ $1.7^{+0.2}_{-0.4}$
ASASSN-19bt $1.6^{+1.2}_{-0.6}$ $0.89^{+0.12}_{-0.14}$ $6.80^{+0.05}_{-0.11}$ $20.79^{+0.03}_{-0.10}$ $8^{+3}_{-4}$ $7^{+1.0}_{-0.1} 0.0$$009^{+0.004}_{-0.005}$ $9.5^{+0.9}_{-1.2}$ $0.4^{+0.2}_{-0.1}$ $2.5^{+0.4}_{-0.2}$
AT2019dsg $0.1^{+2.5}_{-0.1}$ $0.89^{+0.18}_{-0.13}$ $7.07^{+0.14}_{-0.17}$ $19^{+1}_{-2}$ $0.6^{+0.2}_{-0.1}$ $5^{+0.3}_{-0.2} 0.0$$08^{+0.13}_{-0.05}$ $36^{+10}_{-12}$ $0.9^{+0.7}_{-0.3}$ $7.7^{+1.7}_{-2.4}$
AT2019ehz $0.2^{+1.3}_{-0.2}$ $0.46^{+0.10}_{-0.07}$ $6.69^{+0.04}_{-0.04}$ $20.92^{+0.04}_{-0.05}$ $2.7^{+0.5}_{-0.4}$ $0.14^{+0.02}_{-0.03}$ $0.28^{+0.07}_{-0.09}$ $0^{+1}_{-2} 1.1$$1^{+0.5}_{-0.4}$ $3.6^{+0.7}_{-0.5}$
AT2019eve $0.04^{+0.67}_{-0.03}$ $0.14^{+0.13}_{-0.06}$ $6.4^{+0.2}_{-0.2}$ $19^{+2}_{-2}$ $10^{+24}_{-5}$ $5^{+0.4}_{-0.2} 0.0$$02^{+0.04}_{-0.02}$ $16^{+4}_{-4}$ $0.03^{+0.03}_{-0.01}$ $0.10^{+0.01}_{-0.04}$
AT2019qiz $5^{+1}_{-4}$ $0.40^{+0.07}_{-0.26}$ $6.19^{+0.07}_{-0.16}$ $20.6^{+0.2}_{-0.7}$ $3.6^{+2.9}_{-0.5}$ $9^{+2.6}_{-0.2} 0.0$$004^{+0.002}_{-0.002}$ $7^{+1}_{-3}$ $0.22^{+0.09}_{-0.08}$ $0.3^{+0.1}_{-0.2}$
AT2019teq $6^{+13}_{-16} 1.1$$1^{+0.4}_{-0.3}$ $5.8^{+0.3}_{-0.2}$ $20.93^{+0.13}_{-0.12}$ $4^{+3}_{-2}$ $7^{+0.6}_{-0.4} 0.0$$006^{+0.012}_{-0.004}$ $21^{+8}_{-18}$ $5^{+8}_{-3}$ $0.6^{+0.1}_{-0.5}$
AT2020pj $0.07^{+0.94}_{-0.06}$ $0.8^{+0.2}_{-0.2}$ $6.5^{+0.2}_{-0.2}$ $21.21^{+0.06}_{-0.07}$ $5^{+7}_{-2}$ $4^{+0.2}_{-0.2} 0.0$$010^{+0.014}_{-0.006}$ $17^{+5}_{-9}$ $0.5^{+1.3}_{-0.3}$ $1.4^{+0.6}_{-0.7}$
AT2020neh $0.05^{+0.40}_{-0.05}$ $0.3^{+0.4}_{-0.1}$ $6.18^{+0.14}_{-0.14}$ $18^{+2}_{-1}$ $10^{+300}_{-7}$ $6^{+0.4}_{-0.2} 0.0$$02^{+0.10}_{-0.02}$ $8^{+3}_{-5}$ $0.8^{+3.7}_{-0.7}$ $0.6^{+0.4}_{-1.2}$
AT2020nov $1.0^{+3.1}_{-0.8}$ $0.99^{+0.04}_{-0.08}$ $7.26^{+0.03}_{-0.07}$ $18^{+2}_{-2}$ $3.3^{+0.6}_{-0.9}$ $31^{+0.13}_{-0.04} 0.0$$010^{+0.002}_{-0.002}$ $16^{+2}_{-2}$ $0.034^{+0.050}_{-0.004}$ $0.85^{+0.05}_{-0.58}$
AT2020wey $0.1^{+6.1}_{-0.1}$ $0.8^{+0.2}_{-0.2}$ $6.59^{+0.11}_{-0.14}$ $20.74^{+0.08}_{-0.14}$ $24^{+84}_{-19}$ $6^{+0.3}_{-0.1} 0.0$$006^{+0.003}_{-0.002}$ $19^{+3}_{-3}$ $0.4^{+0.5}_{-0.2}$ $1.4^{+0.5}_{-0.6}$
Early vs. Late No Detection + Host vs. Late No Detection + Host vs. Early
$M_{\rm BH,host}$ 0.050 0.049 0.227
$\log M_{\rm BH,mosfit}$ >0.25>0.25 0.199 >0.25>0.25
$t_{\rm viscous}$ 0.188 >0.25>0.25 >0.25>0.25
$b$ >0.25>0.25 0.196 >0.25>0.25
$\log n_{\rm H,host}$ 0.132 0.074 >0.25>0.25
$R_{\rm ph,0}$ 0.026 0.056 >0.25>0.25
$M_{*}$ >0.25>0.25 >0.25>0.25 >0.25>0.25
$\epsilon$ >0.25>0.25 >0.25>0.25 >0.25>0.25
$l_{\rm photo}$ >0.25>0.25 >0.25>0.25 >0.25>0.25
$t_{\rm rise}$ >0.25>0.25 0.151 >0.25>0.25
$\dot{M}_{\rm max}$ >0.25>0.25 >0.25>0.25 >0.25>0.25
$L_{\rm max}$ >0.25>0.25 >0.25>0.25 >0.25>0.25
@@
Figure 3: Same as Figure 2 for the remaining TDEs in our sample with a delayed radio emission episode and converged MOSFiT models. For six of these TDEs, C24 provide a rough estimate of the outflow launch time (dashed vertical red lines). For AT2019dsg, the dashed line indicates the possible launch time of the second radio flare. For PS16dtm, the open circle denotes a marginal possible detection discovered in our reanalysis of the data (Section II.1). C24 do not estimate a launch time for AT2020neh due to its faintness.
@@
III.3 MOSFiT modeling: accretion rate
@@
In Figures 2-4 we show the $\dot{M}$ curves together with the radio and X-ray light curves for all TDEs from our sample with unambiguously transient radio emission and converged MOSFiT models777i.e., all TDEs with radio type “Prompt” or “Delayed” in Table 1 except for AT2018fyk, whose optical emission cannot be well-fit by MOSFiT. Further late-time radio monitoring of the TDEs with radio types Host or None is ongoing and will be presented in future work.. Specifically, Figure 2 shows TDEs for which we have a reasonably well-constrained estimate of the outflow launch time [^8] ASASSN-14ae, AT2018hco, AT2018hyz, and AT2019eve have multiple radio SEDs at late times, so it is possible to use the evolution of the derived physical size of the radio-emitting region to obtain a rough estimate of the time at which their delayed outflows were launched (Cendes et al. 2022a and C24). Similarly, Horesh et al. (2021b) model the multi-frequency radio light curves of iPTF16fnl to constrain the likely radio outflow launch time to $\sim 56-153$ days post-discovery. Multi-wavelength observations of ASASSN-15oi suggest that an outflow launched at the time of the soft X-ray peak may provide a good explanation for its second radio flare (Horesh et al., 2021a; Hajela et al., 2025)., Figure 3 shows TDEs with delayed radio emission where we have at most an approximate constraint on the outflow launch time from C24999plus AT2019dsg, for which we show the approximate onset time of its secondary delayed peak from C24, and AT2020neh, which is too faint for an accurate launch estimate., and Figure 4 shows the remaining TDEs with prompt radio emission. We normalize $\dot{M}$ to $\dot{M}_{\rm Edd}$, where $\dot{M}_{\rm Edd}$ is the accretion rate required to power emission equal to the Eddington luminosity of the SMBH assuming a radiative efficiency $\epsilon$; $L_{\rm Edd}=\epsilon\dot{M}_{\rm Edd}c^{2}$. We use the radiative efficiency from MOSFiT for each TDE and our two independent estimates of the SMBH mass from Section III.1. We find that most TDEs in our sample achieve a peak $\dot{M}$ that approaches or mildly exceeds $\dot{M}_{\rm Edd}$. The peak radio luminosity of a TDE does not appear to correlate with the maximum $\dot{M}/\dot{M}_{\rm Edd}$ achieved onto its SMBH.
@@
We next investigate if the onset time of the radio emission in our TDEs correlates with any specific $\dot{M}/\dot{M_{\rm Edd}}$ ratio, beginning with the TDEs in Figure 2 where we have the best constraints. We find that three of these TDEs (AT2018hco, ASASSN-15oi, and iPTF16fnl) have outflows that were plausibly launched when the accretion rate was a few percent of $\dot{M}_{\rm Edd}$, although for the latter two objects we are limited by the systematic uncertainties on the SMBH masses. ASASSN-15oi’s and iPTF16fnl’s radio emission may also have been launched earlier, possibly when $\dot{M}$ was still super-Eddington. [^10] This is particularly likely for ASASSN-15oi, whose multi-wavelength observations suggest that accretion in this TDE may be delayed a few hundred days relative to our MOSFiT prediction (Hajela et al., 2025). AT2018hyz and AT2019eve’s radio outflows do not appear until several hundred days after the $0.03\,\dot{M}/\dot{M}_{\rm Edd}$ transition, implying that either accretion is delayed relative to the MOSFiT prediction for these two events, or that some other process is responsible for their radio brightenings. [^11] The launch time from Cendes et al. (2022a) assumes that AT2018hyz’s radio outflow is non-relativistic; if its radio emission is instead due to a relativistic off-axis jet, then a much earlier launch time coinciding with the predicted period of super-Eddington accretion is possible (Matsumoto and Piran, 2023; Sfaradi et al., 2024). Finally, ASASSN-14ae’s very late outflow launch time is difficult to reconcile with either state transition, assuming the outflow launch time and the mass of its SMBH are both correct. This may suggest that either accretion disk formation is significantly delayed in this TDE relative to the prediction from MOSFiT, or some other unrelated process is required to explain its radio rise.
@@
Although the onset time of radio emission is less clear for the TDEs in Figure 3, we can roughly approximate the launch time of each outflow based on the radio light curve (C24; dashed red lines in Figure 3). [^12] We note that our new tentative detection of PS16dtm (Section II.1; open circle in Figure 3) pre-dates the launch time suggested by C24 for this TDE; it is therefore possible that its radio outflow was launched significantly earlier ($\sim 1$ year post-discovery). The systematic uncertainties on the SMBH masses and, in some cases, the sparse temporal sampling of the radio light curves preclude a definitive association of the radio emission with a particular accretion rate for these TDEs. It is nevertheless allowed within the uncertainties that the delayed radio emission in all of these events (including the secondary peak in the light curve of AT2019dsg) is driven by a state change in the accretion flow associated with $\dot{M}$ declining to $\lesssim 0.03\,\dot{M}_{\rm Edd}$. [^13] PS16dtm and AT2018zr’s outflows may also be launched at significantly lower accretion rates, depending on which black hole mass measurement is preferred. Again, in most cases, peak accretion onto the SMBH would have to be delayed by several hundred days or more relative to the MOSFiT predictions to associate the outflow launch times with super-Eddington accretion instead. [^14] PS16dtm and AT2019ehz are the two possible exceptions, if their host galaxy-derived SMBH masses rather than their MOSFiT-derived SMBH masses are assumed to be correct.
@@
III.4 X-ray light curves and the Fundamental Plane
@@
The X-ray properties of our TDEs may also connect to the instantaneous accretion rate onto the SMBH (e.g., Metzger 2022; Sarin and Metzger 2024). Together with our SMBH mass measurements (Section III.1), we can use our radio and X-ray light curves to track the evolution of individual TDEs in the $L_{X}$-$L_{R}$ phase space over thousands of days and compare the properties of our sample to the Fundamental Plane of black hole activity. The Fundamental Plane is a well-established correlation between radio luminosity, X-ray luminosity, and black hole mass that holds across many orders of magnitude for black holes accreting at low rates (Merloni et al., 2003; Falcke et al., 2004). If delayed radio emission is produced in TDEs after the accretion rate has dropped to similarly low levels, we may expect these TDEs to obey the same relation at late times.
We show our TDE sample on the Fundamental Plane in Figure 5, using the SMBH masses derived from the host galaxy properties [^15] Using $M_{\rm BH,host}$ rather than $M_{\rm BH,mosfit}$ allows us to include DES14C1kia and AT2018fyk, which do not have converged MOSFiT models. (Table 2). The original sample of AGN from Merloni et al. (2003) are also shown for comparison. As this relation is calibrated using radio observations collected at 5 GHz, we restrict ourselves to radio data taken at similar frequencies ($\nu<10$ GHz). We furthermore only show TDEs for which quasi-simultaneous radio and X-ray data are available, i.e., observations for which $\delta t/t<0.25$. For TDEs with high-cadence observations, if multiple radio/X-ray data point pairs fall within this tolerance, we select the data pairs with the smallest separation in time and frequency space, such that each available data point is used at most once in the figure. 29 of our 31 TDEs have at least one radio/X-ray data point pair satisfying these restrictions and are included in Figure 5. We highlight six TDEs in the figure legend that illustrate the diverse range of radio and X-ray behaviors seen in our sample. Among the TDEs with sufficiently good radio and X-ray coverage, we see that all of our prompt radio TDEs and many of our delayed radio TDEs are consistent with the Fundamental Plane (22 of 29 events). Conversely some TDEs move away from the Plane at late times. The existing data show clear movement away from the plane for two TDEs (ASASSN-15oi and AT2018hyz). Several other TDEs may also exhibit this trend (most suggestively ASASSN-19bt, also potentially PS16dtm and AT2019ehz), but the limitations of the current data preclude a definitive analysis.
IV Discussion
We now attempt to synthesize our multi-wavelength analyses into a physical model, or models, to explain delayed radio emission in TDEs. Any model must be able to explain our key observational findings from Section III, which we briefly summarize here: